# Global identification of dynamic panel models with interactive effects

### v1

@@ 4.2 Intermediate results
The determinant of every diagonal exclusion minor of dimension $k$ of $\mathbf{O}$ can be written as the sum of the determinant of the corresponding diagonal exclusion minor for $\mathbf{\Omega}$ and a term expressed as $(\alpha-\widetilde{\alpha})\widetilde{J}\left(\widetilde{\alpha},\theta\right)$
In mathematical terms, what we will show is that given any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$, say $\mathbf{M}^{\mathbf{O},k}_{R,C}$, we can write its determinant as the sum of the determinant of the diagonal exclusion minor of $\mathbf{\Omega}$ associated with the same rows and columns, say $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$, and a term that can be written as $(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ (10)
where $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a scalar that is a function of the true structural parameters $\theta$ and $\widetilde{\alpha}$. Note that the value of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ will depend on which columns and rows were taken from the original matrix so in general, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)\neq\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)$ for any other valid set of rows and columns $R^{\prime}$ and $C^{\prime}$. To keep notation as simple as possible, we sometimes use $\widetilde{J}_{R,C}$ instead of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ to represent the same object.
Proof: We can prove Proposition 1 by induction.
@@
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ (12)
@@
Let $\mathbf{M}^{\mathbf{O},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{O}$ of dimension $k$ with rows $R=\{r_{1},r_{2},...,r_{k}\}$ and columns $C=\{c_{1},c_{2},...,c_{k}\}$ satisfying the usual conditions. Before proceeding, we need to to establish a correspondence between the indexes of rows $R$ and columns $C$ in the original matrices (either $\mathbf{O}$ or $\mathbf{\Omega}$) and their new indexes (i.e. $\{1,2,...,k\}$) in the $k\times k$ minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. Without loss of generality, let $(i,j)$ denote the position of the $i$-th row in $R$ and the $j$-th column in $C$ within the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. This establishes that the $(i,j)$-th element of the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ are respectively given by $O_{r_{i},c_{j}}$ and $\Omega_{r_{i},c_{j}}$.
@@
where $k$ and $j$ represent the indexes in minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$ of row $r_{k}$ and columns $c_{j}$ of the matrix $\mathbf{O}$. The scalar $O_{r_{k},c_{j}}$ is the $(k,j)$-th element of the minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$; the matrix $\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}$ is the minor of matrix $\mathbf{O}$ with dimension $k-1$ that includes rows $R-r_{k}$ and columns $C-c_{j}$.
@@
Next, define $\widetilde{J}_{R,C}$ as
$\widetilde{J}_{R,C}=\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r,C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]$ (14)
@@
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}$
$\Box$
@@
Consider any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$. Then, $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta)$ is a random polynomial in $\widetilde{\alpha}$ of order in between $\max\{0,k-\bar{r}\}$ and $12k-$ 1.
@@
Proof: First, note from (14) that $\widetilde{J}_{R,C}$ is recursively defined, so once we know $\widetilde{J}_{R,C}$ for $|R|=|C|=1$ we can compute $\widetilde{J}_{R,C}$ for any arbitrary set of rows and columns.
@@
When $k=1$, we have shown that $\widetilde{J}_{R,C}=J_{r,c}$. Now, recall the definition of $\mathbf{J}$ in Claim 2. $\mathbf{\Omega}$ and $\mathbf{L}$ do not depend on $\widetilde{\alpha}$, so $J_{r,c}$ is a polynomial of order 1 in $\widetilde{\alpha}$ if $r>\bar{r}+1$ and $c\neq 1$ or $c>\bar{r}+1$ and $r\neq 1$, and it is a polynomial of order 0 in $\widetilde{\alpha}$ if $r$ or $c$ are equal to 1 (see the discussion about $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ in Corollary 2). Intermediate cases will depend on the values of the other parameters, but it will be a polynomial of either order 0 or 1, depending on whether $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}$ is equal or different from zero. Given the recursive structure of $\widetilde{J}_{R,C}$ we can conclude that for any value of $k$, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a polynomial in $\widetilde{\alpha}$.
@@
Note that the coefficients of this polynomial depend on the true data-generating process parameters $\theta$, and that by Assumptions 5 and 6, $\mathbf{F}$ is drawn from a continuous probability distribution. Then, we can conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ and that the randomness in the coefficients is given by the factor realizations $\mathbf{F}$.
@@
Having this in mind, we first show that the proposed upper bound for the order of the polynomial holds. Then, we show that the lower bound holds as well.
@@
1. Proof for the upper bound: We will show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial of at most order $12k-$ 1 when $|R|=|C|=k$. The proof will be done by induction.
@@
Base case $k=1$: We have shown that when $k=1$ $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $11=2k-$ 1, so the step base holds.
@@
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12(k-1)-$ 1.
Inductive step: Recall the recursive definition of $\widetilde{J}_{R,C}$ derived in (14). By the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12(k-1)-$ 1, $(\alpha-\widetilde{\alpha})$ is a polynomial of order 1 in $\widetilde{\alpha}$, and $J_{r_{k},c_{j}}$ is also a random polynomial in $\widetilde{\alpha}$ of at most order 1 (already discussed). Therefore, the term $(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12k-$ 1. Since the first two terms are of lower order ($\mathbf{\Omega}$ does not depend on $\widetilde{\alpha}$), we conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most order $12k-$ 1. $\square$
@@
2. Proof for the lower bound: First, note that when $k\leq\bar{r}$ our statement trivially holds because $\max\{0,k-\bar{r}\}=0$ and $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ is a polynomial in $\widetilde{\alpha}$. Then, we are left to show that our statement holds for $k>\bar{r}$. We prove this by induction:
Base case $k=\bar{r}+1$: If $k=\bar{r}+1$, then it must be the case that $\exists\ c_{j}\in C$ such that $c_{j}>\bar{r}$. Thus, by Corollary 2 we know there exists a $c_{j}$ such that $J_{r_{k},c_{j}}\neq 0$ with probability 1. As by the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial in $\widetilde{\alpha}$ of at least order 00 and $(\alpha-\widetilde{\alpha})$ is of order 1, then $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ should be at least of order $1=k-\bar{r}=\max\{k-\bar{r},0\}$, as desired.
@@
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$, with $k-1>\bar{r}$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}$ is a random polynomial in $\widetilde{\alpha}\left(\widetilde{\alpha},\theta\right)$ of at least order $\max\{0,k-1-\bar{r}\}=k-1-\bar{r}$.
Inductive step: For any $k>\bar{r}$ we know that there exists a $c_{j}$ such that $J_{r_{k},c_{j}}\neq 0$ with probability 1 (Corollary 2). By the inductive hypothesis $\widetilde{J}_{R-r_{k},C-c_{j}}$ is a random polynomial of at least order $k-1-\bar{r}$, $(\alpha-\widetilde{\alpha})$ is of order 1, which implies that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)$ should be a random polynomial of at least order $k-\bar{r}=\max\{0,k-\bar{r}\}$, as desired. $\square$
@@
where $\mathbf{D}$ is a diagonal matrix. Then, one can see that all the non-diagonal terms of $\mathbf{\Omega}$ are equal to the non-diagonal terms of $\mathbf{F\Psi F^{\prime}}$. Thus, as all the terms are equal it is obvious that for every diagonal exclusion $k\times k$ minor we have that
@@
Proof: This is a direct implication from the previous Claim. First, note that from Claim 3, the determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ is equal to the determinant of the corresponding non-diagonal minor of $\mathbf{F\Psi F^{\prime}}$
@@ 4.3 Proof of Theorem 1
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$
@@
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=0$ (15)
By Assumption 3 we know that $\mathbf{O}$ has at least $3$ different diagonal exclusion minors. Then, consider any other diagonal exclusion minor of $\mathbf{O}$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}$. Using the same logic, we can see that:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$
Then, it must be the case that either $\widetilde{\alpha}=\alpha$ or $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ in order for (5) to hold (Claim 1). However, we will show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta\right)=0$ is only possible in a set of probability zero.
@@
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\left(\alpha-\widetilde{\alpha}\right)\Big{(}\left[\Psi f_{3}(d_{2}-\alpha d_{1}f_{2})\right]\widetilde{\alpha}-\Psi f_{4}(d_{2}-\alpha d_{1}f_{2})\Big{)}=0$ (16)
where in this case $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta\right)$ is a random polynomial in $\widetilde{\alpha}$ of order 1. Note that one of the following four alternatives must be true in order for (16) to be satisfied:
@@
Cases 1 and 2 correspond to the case in which $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta\right)$ is the 0-polynomial, so that all the coefficients are trivially equal to 0. Case 3 corresponds to the case in which we can found a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta)=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the 0-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 4 corresponds to the case where global identification is achieved.
We will show that Cases 1-3 are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 4 must hold so $\alpha$ is almost surely globally identified.
@@
From Assumptions 5 and 6, we know that $d_{1},d_{2},\alpha$ are fixed population parameters, whereas $f_{2}$ is drawn from a continuous probability distribution. Thus, we can invoke Lemma 2 and ensure that $P\left(f_{2}=\frac{d_{2}}{d_{1}\alpha}\right)=0$. Then, Case 2 can only occur in a set of probability 0.
@@
Setting $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$ so that $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta)=0$ leaves us with no other free parameter. Therefore, in order for Case 3 to be possible, it must be the case that this same $\widetilde{\alpha}^{\star}$ simultaneously make all the $\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star},\theta)$ equal to 0 for any other valid set of rows and columns (in particular, it requires $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha}^{\star},\theta)=0$). To see why this will not be the case, consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
@@
$\det\left(\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=0$
Algebra calculation shows that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)$ is a polynomial of order 2 in $\widetilde{\alpha}$:
$\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta\right)=a\widetilde{\alpha}^{2}+b\widetilde{\alpha}+c$ (17)

### v2

@@ 4.2 Intermediate results
The determinant of every diagonal exclusion minor of dimension $k$ of $\mathbf{O}$ can be written as the sum of the determinant of the corresponding diagonal exclusion minor for $\mathbf{\Omega}$ and a term expressed as $(\alpha-\widetilde{\alpha})\widetilde{J}\left(\widetilde{\alpha},\theta^{0}\right)$
In mathematical terms, what we show is that given any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$, say $\mathbf{M}^{\mathbf{O},k}_{R,C}$, we can write its determinant as the sum of the determinant of the diagonal exclusion minor of $\mathbf{\Omega}$ associated with the same rows and columns, say $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$, and a term that can be written as $(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ (10)
where $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a scalar that is a function of the true structural parameters $\theta^{0}$ and $\widetilde{\alpha}$. Note that the value of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ will depend on which columns and rows were taken from the original matrix so in general, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)\neq\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)$ for any other valid set of rows and columns $R^{\prime}$ and $C^{\prime}$. To keep notation as simple as possible, we sometimes use $\widetilde{J}_{R,C}$ instead of $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ to represent the same object.
Proof.
We can prove Proposition 1 by induction.
@@
$\det\left(\mathbf{M}^{\mathbf{O},k-1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ (12)
@@
Let $\mathbf{M}^{\mathbf{O},k}_{R,C}$ be a diagonal exclusion minor of $\mathbf{O}$ of dimension $k$ with rows $R=\{r_{1},r_{2},...,r_{k}\}$ and columns $C=\{c_{1},c_{2},...,c_{k}\}$ satisfying the usual conditions. Before proceeding, we need to establish a correspondence between the indices of rows $R$ and columns $C$ in the original matrices (either $\mathbf{O}$ or $\mathbf{\Omega}$) and their new indices (i.e. $\{1,2,...,k\}$) in the $k\times k$ minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. Without loss of generality, let $(i,j)$ denote the position of the $i$-th row in $R$ and the $j$-th column in $C$ within the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$. This establishes that the $(i,j)$-th element of the minors $\mathbf{M}^{\mathbf{O},k}_{R,C}$ and $\mathbf{M}^{\mathbf{\Omega},k}_{R,C}$ are respectively given by $O_{r_{i},c_{j}}$ and $\Omega_{r_{i},c_{j}}$.
@@
where $k$ and $j$ represent the indices in minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$ of row $r_{k}$ and columns $c_{j}$ of the matrix $\mathbf{O}$. The scalar $O_{r_{k},c_{j}}$ is the $(k,j)$-th element of the minor $\mathbf{M}^{\mathbf{O},k}_{R,C}$; the matrix $\mathbf{M}^{\mathbf{O},k-1}_{R-r_{k},C-c_{j}}$ is the minor of matrix $\mathbf{O}$ with dimension k1k-1 that includes rows $R-r_{k}$ and columns $C-c_{j}$.
@@
Next, define $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ as
$\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})=\sum_{j=1}^{k}(-1)^{k+j}\left[J_{r_{k},c_{j}}\det\left(\mathbf{M}^{\mathbf{\Omega},k-1}_{R-r_{k},C-c_{j}}\right)+\Omega_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}+(\alpha-\widetilde{\alpha})J_{r_{k},c_{j}}\widetilde{J}_{R-r_{k},C-c_{j}}\right]$ (14)
@@
$\det\left(\mathbf{M}^{\mathbf{O},k}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},k}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$
$\square$
@@
Consider any diagonal exclusion minor of $\mathbf{O}$ of dimension $k$. Then, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree almost surely in between $\max\{0,k-\bar{r}\}$ and 2k12k-1.
Proof.
@@
First, note from (14) that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is recursively defined, so once we know $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ for $|R|=|C|=1$ we can compute $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ for any arbitrary set of rows and columns.
@@
When $k=1$, we have shown that $\widetilde{J}_{R,C}=J_{r,c}$. Now, recall the definition of $\mathbf{J}$ in Claim 2. $\mathbf{\Omega}$ and $\mathbf{L}$ do not depend on $\widetilde{\alpha}$, so $J_{r,c}$ is a polynomial of degree 1 in $\widetilde{\alpha}$ if $r>\bar{r}+1$ and $c\neq 1$ or $c>\bar{r}+1$ and $r\neq 1$, and it is a polynomial of degree 0 in $\widetilde{\alpha}$ if $r$ or $c$ are equal to 1 (see the discussion about $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ in Corollary 2). Intermediate cases will depend on the values of the other parameters, but it will be a polynomial of either degree 0 or 1, depending on whether $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}$ is equal or different from zero. Given the recursive structure of $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ we can conclude that for any value of $k$, $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial in $\widetilde{\alpha}$.
@@
Note that the coefficients of this polynomial depend on the true data-generating process parameters $\theta^{0}$, and that by Assumptions A6 and A7, $\mathbf{F}$ is drawn from a continuous probability distribution. Then, we can conclude that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ and that the randomness in the coefficients is given by the factor realizations $\mathbf{F}$.
@@
Having this in mind, we first show that the proposed upper bound for the degree of the polynomial holds. Then, we show that the lower bound holds as well.
@@
1. Proof for the upper bound: We show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial of at most degree 2k12k-1 when $|R|=|C|=k$. The proof will be done by induction.
@@
Base case $k=1$: We have shown that when $k=1$ $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most degree 1=2k11=2k-1, so the step base holds.
@@
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of at most degree 2(k1)12(k-1)-1.
[…]
@@
2. Proof for the lower bound: First, note that when $k\leq\bar{r}$ our statement trivially holds because $\max\{0,k-\bar{r}\}=0$ and $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial in $\widetilde{\alpha}$. Then, we are left to show that our statement holds for $k>\bar{r}$. We prove this by induction:
[…]
@@
Inductive hypothesis: Assume our proposition holds for diagonal exclusion minors of dimension $(k-1)\times(k-1)$, with $k-1>\bar{r}$. This implies that for any set of valid $R$ and $C$ we have that $\widetilde{J}_{R,C}(\widetilde{\alpha},\theta^{0})$ is a random polynomial in $\widetilde{\alpha}$ of at least degree $\max\{0,k-1-\bar{r}\}=k-1-\bar{r}$.
[…]
@@
where $\mathbf{D}$ is a diagonal matrix. Then, one can see that all the non-diagonal terms of $\mathbf{\Omega}$ are equal to the non-diagonal terms of $\mathbf{F\Psi F^{\prime}}$. Thus, as all the terms are equal it is evident that for every diagonal exclusion $k\times k$ minor we have that
@@
Proof.
This is a direct implication from the previous Claim. First, note that from Claim 3, the determinant of every diagonal exclusion minor of $\mathbf{\Omega}$ is equal to the determinant of the corresponding non-diagonal minor of $\mathbf{F\Psi F^{\prime}}$
@@ 4.3 Proof of Theorem 1
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=\det\left(\mathbf{M}^{\mathbf{\Omega},\bar{r}+1}_{R,C}\right)+(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$
@@
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R,C}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=0$ (15)
By Assumption A3 we know that $\mathbf{O}$ has at least $2$ different diagonal exclusion minors. [^4] In particular, when $T=2(\bar{r}+1)$ and $\bar{r}>0$ there are at least $\frac{\binom{2(\bar{r}+1)}{\bar{r}+1}}{2}\geq 2$ diagonal exclusion minors Then, consider any other diagonal exclusion minor of $\mathbf{O}$, say $\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}$. Using the same logic, we can see that:
$\det\left(\mathbf{M}^{\mathbf{O},\bar{r}+1}_{R^{\prime},C^{\prime}}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Then, it must be the case that either $\widetilde{\alpha}=\alpha$ or $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ in order for (5) to hold (Claim 1). However, we show that $\widetilde{J}_{R,C}\left(\widetilde{\alpha},\theta^{0}\right)=\widetilde{J}_{R^{\prime},C^{\prime}}\left(\widetilde{\alpha},\theta^{0}\right)=0$ is only possible in a set of probability zero.
@@
$\det\left(\mathbf{M}_{(1,2),(3,4)}^{\mathbf{O},2}\right)=\left(\alpha-\widetilde{\alpha}\right)\Psi(d_{2}-\alpha d_{1}f_{2})\left(f_{3}\widetilde{\alpha}-f_{4}\right)=0$ (16)
where in this case $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a random polynomial in $\widetilde{\alpha}$ of degree 1. Note that one of the following four alternatives must be true in order for (16) to be satisfied:
@@
Cases 1 and 2 correspond to the case in which $\widetilde{J}_{(1,2),(3,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is the zero polynomial, so that all the coefficients are trivially equal to 0. Case 3 corresponds to the case in which we can find a root $\widetilde{\alpha}^{\star}$ that solves $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$. However, once $\widetilde{\alpha}$ is pinned-down in this way, we have no other free parameter, so the same root must satisfy all the zero-determinant conditions (Claim 1) associated with the remaining diagonal exclusion minors of $\mathbf{O}$. Case 4 corresponds to the case where global identification is achieved.
We show that Cases 1-3 are only possible in a set of factor realizations that occur with probability zero. Thus, with probability 1, Case 4 must hold so $\alpha$ is almost surely globally identified.
@@
From Assumptions A6 and A7, we know that $d_{1},d_{2},\alpha$ are fixed population parameters, whereas $f_{2}$ is drawn from a continuous probability distribution. Thus, we can invoke Lemma 2 and ensure that $P\left(f_{2}=\frac{d_{2}}{d_{1}\alpha}\right)=0$. Then, Case 2 can only occur in a set of probability 0.
@@
Setting $\widetilde{\alpha}^{\star}=\frac{f_{4}}{f_{3}}$ so that $\widetilde{J}_{(1,2),(3,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$ leaves us with no other free parameter. Therefore, in order for Case 3 to be possible, it must be the case that this same $\widetilde{\alpha}^{\star}$ simultaneously make all the $\widetilde{J}_{R^{\prime},C^{\prime}}(\widetilde{\alpha}^{\star},\theta^{0})$ equal to 0 for any other valid set of rows and columns (in particular, it requires $\widetilde{J}_{(2,3),(1,4)}(\widetilde{\alpha}^{\star},\theta^{0})=0$). To see why this will not be the case, consider the following $2\times 2$ diagonal exclusion minor of $\mathbf{O}$:
@@
$\det\left(\mathbf{M}_{(2,3),(1,4)}^{\mathbf{O},2}\right)=(\alpha-\widetilde{\alpha})\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=0$
Algebra calculation shows that $\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)$ is a polynomial of degree 2 in $\widetilde{\alpha}$:
$\widetilde{J}_{(2,3),(1,4)}\left(\widetilde{\alpha},\theta^{0}\right)=a\widetilde{\alpha}^{2}+b\widetilde{\alpha}+c$ (17)